# SMARTER: A Data‑efficient Framework to Improve Toxicity Detection with Explanation via Self-augmenting Large Language Models

### v1

@@ 6.2.1 Cross-model Refinement Results
(a) HateXplain
(b) Latent Hate
(c) Implicit Hate
6.2.2 Stylistic Analysis
From Figure 5, most explanations from the individually trained models exhibit their corresponding style, while cross-model SFT fine-tuning is sufficient to substantially align the outputs with the alternative’s style. With the exception of T5 on HateXplain, the application of self-augmenting DPO using cross-model data generally amplifies the stylistic presence of the base model to a small extent, suggesting that this mechanism induces a mild converging effect on the reasoning style.
Cross-model explanations improve weak categories.
We analyze how cross-model refinement enhances classification by comparing Precision, Recall, and F1-score breakdowns of the base models and their cross-trained counterparts at $K=256$ in Figure 8. Models often improve in categories where they initially struggled but their counterpart performed well, suggesting they can adopt effective reasoning patterns. However, performance may decline if the counterpart model is weak. This tendency, along with the stylistic transference shown in Figure 5, aligns with findings on the benefits of reasoning diversity by incorporating outputs from other models Hao et al. ; Chenglin et al. (2024).
@@ 7 Discussion and Conclusion
T5’s superior F1-scores compared to individually trained models (Figure 4) suggest added value in leveraging competing models. Practitioners can train multiple LLMs using the SFT+DPO framework across data thresholds, then select a high-quality checkpoint (e.g., $K=128$ from Section 5.2.1) for further finetuning with outputs from other models. Though additional performance gain may not be guaranteed, cross-model refinement is still a venue worthy of exploration particularly when resources are limited.
@@
The differing post-training behaviors of T5 and Llama—despite being trained on the same seed explanations—may stem from their distinct architectures (encoder-decoder vs. encoder-only). We hypothesize, based on prior work, that the stronger model (Llama) enhances reasoning diversity that T5 can leverage. While further study is needed for a definitive explanation, our findings suggest practitioners may benefit from using architecturally diverse models within this framework.
@@ 8 Limitations
While our framework aims to improve performance and versatility in content moderation, it does not alleviate the fundamental sensitive nature of this task. Content moderation could be misappropriated to suppress free speech or harm marginalized groups Dias Oliva (2020); Kozyreva et al. (2023). Other works have noted that LLMs exhibit bias behaviors in certain applications Nguyen et al. (2023); Gu et al. (2025). We urge researchers and practitioners to maintain vigilance when integrating our framework with human supervision to ensure ethic standards Lai et al. (2022); Cao et al. (2024).
@@ Appendix B Data Pre-processing
The datasets used in this work are open-source. We peruse them as research artifacts accordingly to their license. To further mitigate risks of privacy, we anonymize posts by replacing replacing user handles with the string <user> and remove URLs if they appear in the context of the posts.
@@ Appendix C Obtaining Explanations for In-domain Data
Appendix D Cross-model Analysis
@@
Appendix E Technical Details
E.1 Hardware
All experiments are carried out on a maximum of 2 Nvidia RTX A6000 GPUs. We download the models from their Huggingface [^3] https://huggingface.co/ checkpoints. Finetuning is implemnted with LoRA techniques Hu et al. (2021) via the TRL library  [^4] https://huggingface.co/docs/trl/en/index. The LoRA configurations for all settings are:
@@ E.1 Hardware
E.2 Preference Optimziation
@@
We primarily use the implementation of the DPOTrainer and KTOTrainer classes via the TRL library hosted on the HugingFace platform. For the DPOTrainer, we set $\beta=0.1$, and the loss to be default sigmoid loss. For the KTOTrainer, we similarly set the $\beta=0.1$ and use the default KTO loss option.
E.3 Hyperparameter Tuning
@@ 64 DPO 3 1e-04 3 1e-06
Appendix F Examples of Explanations

### v2

@@ 6.2.2 Cross-model Classification Results
6.2.3 Cross-model Style Analysis
Table 3: NLI-based consistency between explanations, predicted labels, and label definitions across datasets. Both DPO and cross-refined (XMOD) models show high entailment rates (>96%) with marginal percentages of contradictions (Contra.), neutral or undefined (Undef.) edge cases. Categories with all 0 values are omitted.
Label Consistency (%) Definition Consistency (%)
Dataset Model Train Entail Contra. Undef. Entail Contra. Neutral
HateXplain T5 DPO 99.2 0.8 0.0 98.2 1.5 0.3
XMOD 96.7 1.5 1.8 99.0 0.9 0.2
Llama DPO 97.5 1.2 1.3 97.2 2.6 0.2
XMOD 96.4 3.6 0.0 96.7 3.2 0.1
Latent Hate T5 DPO 98.8 1.2 0.0 97.6 2.4 0.0
XMOD 97.4 1.9 0.7 97.6 2.3 0.1
Llama DPO 97.3 2.7 0.0 97.3 2.3 0.3
XMOD 96.8 3.2 0.0 97.5 2.5 0.0
Implicit Hate T5 DPO 99.6 0.4 0.0 99.0 1.0 0.0
XMOD 98.3 1.4 0.4 97.5 2.0 0.5
Llama DPO 98.1 0.4 1.5 97.5 1.8 0.8
XMOD 96.9 2.8 0.1 96.0 3.6 0.4
We additionally train a style classifier based on the BERT architecture using explanations generated individually by T5 and Llama of the SFT+DPO variant at $K=128$ Devlin et al. (2019). A classification head is added to BERT, which leverages the [CLS] token from each explanation’s embedding to predict whether it was produced by T5 or LLaMA. Technical details are reported in Appendix F.3.
From Figure 11, most explanations from the individually trained models exhibit their corresponding style, while cross-model SFT is sufficient to substantially align the outputs with the alternative’s style. With the exception of T5 on HateXplain, the application of self-augmenting DPO using cross-model data generally amplifies the stylistic presence of the base model to a small extent, suggesting that this mechanism induces a mild converging effect on the reasoning style.
Cross-model explanations may improve weak categories in weaker models.
We compare Precision, Recall, and F1-score of base and cross-trained models at $K=256$ (Figure 10) to assess cross-model refinement. Improvements often occur in categories where the paired model excels, indicating transferable reasoning patterns, while declines appear when the counterpart is weak. This asymmetry underscores our recommendation in line 19 of Algorithm 1 to validate and select the final model using a robust performance metric.
7 Analysis of Explanations
We perform Natural Language Inference (NLI) evaluation to ascertain whether the LLMs’ explanations remain internally consistent with their corresponding labels and definitions throughout the augmentation pipeline, and then anatomize the stylistic changes between cross-model refinement.
7.1 Analysis of Consistency
Explanation-Label Consistency
We prompt GPT-4o-mini with the template shown in Figure 8 to judge whether the offered explanations is logically consistent with the predicted label for each sample across 3 test sets via the 4 categories: Entail, Contradict, Neutral and Undefined (noise cases where either the predicted label or the explanation is missing). An author’s independent annotation of 150 random samples from this pool under the same instruction achieves nearly unanimous agreement with GPT-4o-mini (Cohen’s $\kappa=0.92$).
In Table 3, Entail values exceeding 96% indicate that explanations are largely consistent with their respective labels. However, cross-model training induces a marginal 2-3% increase in Contradiction compared to the DPO augmentation.
Explanation-Definition Consistency
Applying the same method to only the Entail samples from the previous analysis, using the prompt shown in Figure 9 (Cohen’s $\kappa=0.96$), we again observe consistently high entailment rates and a slight increase in contradiction under cross-model training that mirrors the aforementioned trend.
Overall, LLM-generated explanations remain predominantly consistent with both predicted labels and their definitions. Cross-model refinement can enhance classification but may introduce a slight risk of inconsistency, warranting periodic monitoring. Appendix E offers qualitative analysis and insights for further quality checks for deployment.
@@ 8 Discussion and Conclusion
T5’s superior F1-scores compared to individually trained models (Figure 2) suggest added value in leveraging competing models. Practitioners can train multiple LLMs using the SFT+DPO framework across data thresholds, then select a high-quality checkpoint (e.g., $K=128$ from Section 5.2.1) for further finetuning with outputs from other models. Although further gains are not guaranteed, cross-model refinement represents a practical and resource-efficient direction for improving alignment and robustness.
@@
The differing post-training behaviors of T5 and Llama—despite being trained on the same seed explanations—may stem from their distinct architectures (encoder-decoder vs. encoder-only). We hypothesize based on prior work Hao et al. , that the stronger model (Llama) enhances reasoning diversity that T5 can leverage. While further study is needed for a definitive explanation, our findings suggest practitioners may benefit from using architecturally diverse models within this framework.
@@ 9 Limitations
While our framework aims to improve performance and versatility in content moderation, it does not alleviate the fundamental sensitive nature of this task. Content moderation could be misappropriated to suppress free speech or harm marginalized groups Dias Oliva (2020); Kozyreva et al. (2023). Other works have noted that LLMs exhibit bias behaviors in certain applications Nguyen et al. (2023); Gu et al. (2025). We urge researchers and practitioners to maintain vigilance when integrating our framework with human supervision to ensure ethical standards Lai et al. (2022); Cao et al. (2024).
@@ Appendix B Data Pre-processing
The datasets used in this work are open-source. We peruse them as research artifacts accordingly to their license. To further mitigate risks of privacy, we anonymize posts by replacing user handles with the string <user> and remove URLs if they appear in the context of the posts.
@@
Dataset Total Size Split Ratio K val K test Platform Label Space
HateXplain 20,148 60:20:20 50 400 G, T Normal, Offensive, Hate
@@ Appendix C Obtaining Explanations for In-domain Data
Dataset Model Macro F1
HateXplain Llama_OTS 0.52
T5_OTS 0.56
Latent Hate Llama_OTS 0.53
T5_OTS 0.42
Implicit Hate Llama_OTS 0.32
T5_OTS 0.33
Table 5: Comparison of off-the-shelf (OTS) model performance across datasets. LLMs are prompted to produce answers without explanations in this setting as shown in Figure 5.
### Instruction:
By considering their corresponding Definitions, label the following post with only one of these categories: {categories}.
Provide your response in the following format:
LABEL: [text]
### Definitions:
{definitions}
### Post:
{post}
### Response: [model output]
Figure 5: Prompt template for classification tasks without explanation generation. The model is instructed to directly output a label based on the provided definitions.
### Instruction:
Provide your message in the following format:
EXPLANATION: [text]
LABEL: [text]
### Definitions:
Hate: Language that targets [...]
Abusive: Language that is [...]
Normal: Does not contain [...]
### Post:
{post}
### Response: [model output]
@@
### Instruction:
You are given a model’s prediction and its explanation for a social media post.
Your task is to judge whether the explanation is logically consistent with the predicted label.
Possible outputs:
- ENTAIL: The explanation supports and is faithful to the predicted label.
- CONTRADICT: The explanation contradicts the predicted label.
- NEUTRAL: The explanation is unrelated, vague, or does not clearly support the predicted label.
- UNDEFINED: Either the predicted label or the explanation is missing or empty.
Return only one word: ENTAIL, CONTRADICT, NEUTRAL, or UNDEFINED.
### Predicted Label:
{pred_label}
### Model Explanation:
{explanation}
### Response: [model output]
Figure 8: Prompt template for evaluating explanation–label consistency. The model is asked to judge whether a generated explanation is logically consistent with its predicted label.
### Instruction:
You are given a model’s predicted label and its corresponding explanation.
Your task is to judge whether the explanation is logically consistent with the definition of that label.
Possible outputs:
- ENTAIL: The explanation clearly aligns with and supports the definition.
- CONTRADICT: The explanation conflicts with the definition.
- NEUTRAL: The explanation is vague, unrelated, or does not clearly support the definition.
Return only one word: ENTAIL, CONTRADICT, or NEUTRAL.
### Official Definition:
{definition}
### Model Explanation:
{explanation}
### Response: [model output]
Figure 9: Prompt template for evaluating explanation–definition consistency. The model is asked to assess whether a generated explanation logically aligns with the formal definition of the predicted label.
Appendix D Stylistic Analysis
In Figure 11, we show the distribution of Llama vs T5 style pre- and post cross-model refined as recognized by the BERT style classifier.
@@ Appendix D Stylistic Analysis
Appendix E Qualitative Analysis of Explanation
We examine the relatively small amount of non-Entail samples in section 7 to observe general trends in the models’ inconsistency with respect to both the predicted label and definition.
First, T5 models typically exhibit higher consistency in explanation compared to their Llama counterparts as shown in Table 3. Even after cross-model refinement (XMOD), the percentage of Contradiction increases by at most 1% (Implicit Hate). While a conclusive explanation is out of scope, we hypothesize that T5’s encoder-decoder architecture supports higher robustness compared to Llama’s decoder-only.
Second, qualitative analysis reveals different patterns of inconsistency in the explanations offered by these 2 models. For HateXplain and Implicit Hate, Llama’s Contradict-tagged explanations tend to consistently mentions some negative and/or offensive sentiments in the post, yet ultimately yielding the conflicting label Normal (Figure 15). In contrast, T5 Contradict explanations – much fewer in overall number – generally asserts a label at the beginning yet explicitly rejects this label at the end. For the more challenging Implicit Hate, both models’ Contradict explanations typically contain "lost in context" inconsistency, where the justification is grounded in some other labels’ definition rather than the stated one, possibly due to the nuanced complexity of these finer-grained categories.
Our NLI-based consistency check also emerges as a simple yet useful quality check for downstream deployment.
In addition to classification metrics, moderators may opt to implement a secondary lightweight NLI-checker to alert when a threshold of inconsistency is reached.
Appendix F Technical Details
F.1 Hardware
All experiments are carried out on a maximum of 2 Nvidia RTX A6000 GPUs. We download the models from their Huggingface [^3] https://huggingface.co/ checkpoints. Finetuning is implemented with LoRA techniques Hu et al. (2021) via the TRL library  [^4] https://huggingface.co/docs/trl/en/index. The LoRA configurations for all settings are:
@@ F.1 Hardware
F.2 Inference Parameters
All models in our experiments use the following parameters during inference:
  • $\diamond$
    temperature: 0.0
  • $\diamond$
    max_tokens/max_new_tokens:
    • $\diamond$
      512 for classification with explanation
    • $\diamond$
      20 for classification without explanation
The library vLLM [^5] https://docs.vllm.ai/en/latest/ is used coupled with the openAI v1/chatcompletion endpoint to enable fast inference in our experiments. The OpenAI model GPT-5-chat-latest endpoint was accessed on September 12, 2025. We also use the Qwen2.5-32B-Instruct-AWQ version on HuggingFace.
F.3 Preference Optimization
@@
We primarily use the implementation of the DPOTrainer and KTOTrainer classes via the TRL library hosted on the HuggingFace platform. For the DPOTrainer, we set $\beta=0.1$, and the loss to be default sigmoid loss. For the KTOTrainer, we similarly set the $\beta=0.1$ and use the default KTO loss option.
F.4 Hyperparameter Tuning
@@ 64 DPO 3 1e-04 3 1e-06
Appendix G Examples of Explanations
@@
LLAMA - HATEXPLAIN
EXPLANATION: The post targets the group of homosexuals by using the term "queer" in a derogatory manner. The language used attacks and demeans this group based on their sexual orientation, which falls under the definition of hate speech. Therefore, the post is labeled as "Normal."
LABEL: Normal
(a) Llama: The final label ‘Normal’ does not follow from the prior justification.
T5 - LATENT HATE
EXPLANATION: This post should be classified as "Not Hate" because it does not involve any form of hatred, prejudice, or discrimination towards individuals or groups based on their characteristics. The post simply mentions Alex Linder and white nationalism 1.0, without expressing any negative or discriminatory views towards any specific group or individual. Therefore, it does not meet the criteria for being classified as "Not Hate."
LABEL: Not Hate
(b) T5: The explanation starts with an asserted label yet later refutes this category.
LLAMA - IMPLICIT HATE
EXPLANATION: The post uses irony to attack and demean the idea of a wall, sarcastically suggesting that it should also prevent merpeople from the ocean. The use of humor and satire is used to mock the idea of a wall being a barrier to prevent all forms of transportation, including boats, tunnels, and airplanes, as well as the ocean.
LABEL: Threatening
(c) Llama: The explanation discusses factors directly relevant to ‘Irony’, yet the predicted label ‘Threatening’ is unrelated.
Figure 15: Examples of explanations that are judged to contains some forms of inconsistency with respect to the predicted label and/or its definition in our NLI-based analysis.